OpenStudy (anonymous):

What is the arc length of a circle that has an 8-inch radius and a central angle that is 95 degrees? Use 3.14 for π and round your answer to the nearest hundredth.

OpenStudy (anonymous):

@amoodarya @AliceCullen @asmagul @Afrodiddle @AeroSmith @ashontae19 @sweetburger @dan815 @DanJS @ganeshie8 @HELP-ME-PLZ @FoxxJamey @batman19991 @Nicoleegilmoree @Luigi0210 @rebecaadavila @iiAriana @phi @paki

OpenStudy (anonymous):

\[\frac{ central \angle }{360 }=\frac{ arc~length }{ 2\pi r }\]

OpenStudy (batman19991):

A central angle of 60° is 1/6 of the 360° circle. The intercepted arc will have a length of 1/6 of the circumference. circumference = 2πr = 2π(15) = 30π 1/6 circumference = arc = 30π/6 = 5π i dont know if this helps but ya

OpenStudy (anonymous):

These are my choices 760.01 inches 95.08 inches 0.53 inch 13.26 inch

OpenStudy (anonymous):

Which one is it?

OpenStudy (anonymous):

B?

OpenStudy (anonymous):

D. \[\frac{95*3.14*16}{360}=13.2578 \]